OpenStudy (anonymous):

You are going to run at a constant speed of 6.5 miles per hour for 30 minutes. You calculate the distance you will run. What mistake did you make in your calculation? [Use the formula S = dt.] 6.5 • 30 does not equal 195. mi/hr • min does not equal mi. You must convert 30 minutes to hours. d30min • 30 min does not equal d. There are no mistakes. The solution is correct.

OpenStudy (anonymous):

Is there a calculation?

OpenStudy (anonymous):

s= d/t 6.5mi/hr=d/30mi 6.5mi/hr*30mi 195mi=d

OpenStudy (anonymous):

Ok. time needs to be in the same units. You need to convert 30 minutes to 0.5 hours to match miles per hour.

OpenStudy (anonymous):

how am I suppose to do this?

OpenStudy (anonymous):

I'm not sure what you mean. I thought they were just asking you to pick out the mistake.

OpenStudy (anonymous):

I got it, thanks